OpenStudy (rowanmartinfan):

so with a #^# if divided then u + or - the 2 numbers ill say that is a yes so ill try that

OpenStudy (rowanmartinfan):

@jcpd910 \[\frac{ 2^5*8^4 }{ 16 }=\frac{ 2^5*(2^a)^4 }{ 2^4 }= \frac{ 2^5*2^b }{ 2^4 }=2^c\] what is a, b, c

OpenStudy (rowanmartinfan):

@Kainui \[\frac{ 2^5*8^4 }{ 16 }\frac{ 2^5*(2^a)^4}{ 2^4 }=\frac{ 2^5*2^b }{ 2^4 }=2^c\]